# Galaxy Morphologies at Cosmic Noon with JWST: A Foundation for Exploring Gas Transport with Bars and Spiral Arms

### v1

@@
11institutetext: Max-Planck-Institut für extraterrestische Physik (MPE), Giessenbachstraße 1., 85748 Garching, Germany 22institutetext: Max-Planck-Institut für Astrophysik (MPA), Karl-Schwarzschild-Str. 1, 85748, Garching, Germany 33institutetext: Department of Physics, University of Oxford, Clarendon Laboratory, Parks Road, OX1 3PU Oxford, UK 44institutetext: Purple Mountain Observatory, Chinese Academy of Sciences, 10 Yuanhua Road, Nanjing 210023, China 55institutetext: Department of Physics, University of Bath, Claverton Down, Bath BA2 7AY, UK 66institutetext: Research School of Astronomy and Astrophysics, Australian National University, Weston Creek, ACT 2611, Australia 77institutetext: ARC Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), Canberra, ACT 2611, Australia 88institutetext: Kavli Institute for Astronomy and Astrophysics, Peking University, Beijing 100871, People’s Republic of China 99institutetext: Cosmic Dawn Center (DAWN), Denmark 1010institutetext: DTU-Space, Technical University of Denmark, Elektrovej 327, DK2800 Kgs. Lyngby, Denmark $1$institutetext: Department of Physics and Astronomy and PITT PACC, University of Pittsburgh, Pittsburgh, PA 15260, USA 1212institutetext: Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, Padova I-35122, Italy 1313institutetext: School of Physics and Astronomy, Tel Aviv University, Tel Aviv 69978, Israel 1414institutetext: Universidad de Concepción, Víctor Lamas 1290, Barrio Universitario, Concepción, Chile
@@
Juan M. Espejo Salcedo 11 jespejo@mpe.mpg.de    Stavros Pastras 1122    Josef Vácha 33    Claudia Pulsoni 11    Reinhard Genzel 11    N. M. Förster Schreiber 11    Jean-Baptiste Jolly 11    Capucine Barfety 11    Jianhang Chen 11    Giulia Tozzi 11    Daizhong Liu 1144    Lilian Lee 11    Stijn Wuyts 55    Linda Tacconi 11    Ric Davies 11    Hannah Übler 11    Dieter Lutz 11    Emily Wisnioski 6677    Jinyi Shangguan 1188    Minju Lee 991010    Sedona H. Price $1$    Frank Eisenhauer 11    Alvio Renzini 1212    Amit Nestor Shachar 1313    Rodrigo Herrera-Camus 1414
(Received …; accepted …)
How radial flows shape galaxy structure and evolution remains an open question. Internal drivers of such flows, such as bars and spiral arms, known to mediate gas flows in the local Universe, are now observable at high redshift thanks to JWST’s unobscured view. We investigate the morphology of massive star-forming galaxies at $1.30.8<z<$ 1.3 and $2.52.0<z<$ 2.5, epochs marking the peak and decline of cosmic star formation, both well-covered by kinematic surveys. Using JWST/NIRCam imaging, we visually classify 1,451 galaxies, identify non-axisymmetric features, count the number of spiral arms, analyze non-parametric morphological indicators and study the dynamical support of the sample covered by kinematics (10% of the sample) as measured via $v/\sigma$. Disk galaxies dominate the sample (82%), with 48% exhibiting spiral structure and 11% hosting bars. Both fractions decline with redshift, consistent with previous studies. The proportion of two- and three-armed spirals remains largely unchanged across redshift, with roughly two-thirds showing two arms and one-third showing three arms in both bins. Notably, we find a higher incidence of three-armed spirals than reported in the local Universe, suggesting a mild evolution in spiral arm multiplicity. Non-parametric morphological metrics strongly correlate with stellar mass but show no significant redshift evolution. Finally, kinematic analysis reveals a strong correlation between disk morphology and rotational support, with most disks exhibiting $v/\sigma>3$ and median values of $v/\sigma>7$ for spirals and $v/\sigma>5$ for barred galaxies. This study establishes a population-wide framework for linking galaxy morphology and dynamics at cosmic noon, providing a key reference for future studies on the role of detailed structural features in galaxy evolution.
@@ 1 Introduction
A fundamental question in galaxy evolution is how early star-forming galaxies assembled the well-ordered structures seen in the present-day Universe. While early morphological studies suggested that high-redshift galaxies were highly irregular and dynamically unstable, kinematic surveys have since revealed that disk-like rotation is widespread at cosmic noon ($31<z<$ 3) (e.g., Förster Schreiber et al. 2009; Law et al. 2009; Wisnioski et al. 2015; Stott et al. 2016; Simons et al. 2017; Harrison et al. 2017; Übler et al. 2019). Integral field spectroscopy has consistently shown that a large fraction of star-forming galaxies at these epochs exhibit strong rotational support, often with velocity dispersions lower than expected for turbulent, irregular morphologies. These findings challenged the prevailing view based on HST imaging, which primarily identified galaxies with clumpy, chaotic structures (e.g., Abraham et al., 1996; Conselice et al., 2000; Jogee et al., 2004; Lotz et al., 2006; Elmegreen et al., 2007; Sheth et al., 2008).
@@
The arrival of JWST has dramatically transformed our ability to study galaxy structure at high redshift, offering a combination of longer wavelengths, increased depth, and unprecedented resolution. Recent studies leveraging JWST NIRCam imaging consistently reveal a high fraction of disk galaxies at and beyond cosmic noon (e.g., Jacobs et al. 2023; Ferreira et al. 2023; Kartaltepe et al. 2023; Liu et al. 2023; Nelson et al. 2023; Huertas-Company et al. 2024; Lee et al. 2024; Tohill et al. 2024; Pandya et al. 2024). Additionally, some studies provide clear evidence of spiral arms and bars (e.g., Costantin et al. 2023; Ferreira et al. 2023; Guo et al. 2023; Kuhn et al. 2024; Le Conte et al. 2024; Chugunov et al. 2025). Thus, many galaxies that once appeared irregular or clumpy in HST imaging are now recognized as well-ordered disks, with some exhibiting prominent non-axisymmetric features, reaffirming the prevalence of disk from kinematics studies in the pre-JWST time and the now consistent disk fractions inferred from morphology and kinematics-based studies.
Beyond their prevalence, these non-symmetric morphological features are key agents of secular evolution, regulating gas flows and star formation within galaxies. In the local Universe, bars and spiral arms are known to redistribute angular momentum and funnel gas toward galaxy centres, often fueling central star formation and bulge growth. As demonstrated by Yu et al. (2022) with the EDGE-CALIFA survey of local galaxies, both spiral arms and bars efficiently channel gas inward, leading to higher central gas concentrations and elevated star formation rates.
If these mechanisms were already at play at cosmic noon, they could have profound implications for the early buildup of galaxy structure. Recent kinematic studies have uncovered evidence for rapid radial inflows in high-redshift main-sequence disks (e.g., Genzel et al., 2023, Pastras et al. in prep), suggesting that these processes may have been even more pronounced in the past. These findings contradict earlier expectations that non-axisymmetric structures like spiral arms and bars would be short-lived in gas-rich, turbulent environments (e.g., Genzel et al. 2008; Ceverino et al. 2010; Bournaud et al. 2011). Instead, they suggest that such features were not only prevalent but also played a fundamental role in shaping galaxy evolution throughout cosmic time. Given these advancements, establishing a direct connection between morphology and kinematics is crucial for understanding how high-redshift disks evolve and whether non-axisymmetric features are primary drivers of internal gas flows and structural transformation.
@@
Our goal is to establish a robust morphological foundation for interpreting the growing body of kinematic measurements from near-infrared integral field spectroscopy and millimetre/submillimetre observations. Specifically, our sample selection aligns with that of major spectroscopic surveys such as KMOS3D, SINS, KROSS, NOEMA-3D and PHIBSS, thus enabling the connections between the morphological and kinematic analyses. By providing a comprehensive visual classification of galaxy structure (including disks, bars, and spiral arms), this work serves as a key reference for future studies aiming to investigate the connections between morphology, gas transport and secular evolution. Additionally, we complement our visual classifications with quantitative (non-parametric) morphological metrics, leveraging well-established statistical methods to characterize galaxy structure and to systematically link morphology and dynamics at cosmic noon.
In §2, we describe the sample selection, followed by a discussion of the imaging data, its corresponding data reduction methods, and the assembly of the kinematic sample used in this study. §3 presents the overall morphological properties of the sample, detailing the visual classification scheme, general classification results, and the identification of bars and spiral structures. §4 examines the evolution of these morphological features with redshift and compares our findings with previous studies. In §5, we perform a quantitative analysis of galaxy morphology, exploring non-parametric structural indicators and their relationship to the visual classifications. §6 investigates the connection between morphology and kinematics, discussing the implications for galaxy evolution. Finally, §7 summarizes the key findings of this work. Additional details, including discussions of observational limitations, preparations for the morphological analysis, and a description of the methodology used to separate galaxy morphologies based on structural metrics, and comparisons with HST photometry are provided in Appendices A-D.
@@
Figure 1: Star-formation rate (SFR) vs stellar mass $M_{\star}$ for the selected sample, with the Speagle et al. (2014) relations for the mid-points of the redshift bins $z=1.05$ and $z=2.25$.
@@ 2.1 NIRCam imaging
To investigate morphological evolution at cosmic noon, we focus on two epochs corresponding to $1.30.8<z<$ 1.3 and $2.52.0<z<$ 2.5. Since star-forming galaxies dominate the population at these epochs, we apply selection criteria targeting massive star-forming galaxies with star formation rates (SFR) within 0.4 dex of the star formation main sequence (MS), as parametrized by Whitaker et al. (2014), i.e., $\log(\mathrm{SFR}/\mathrm{SFR_{MS}})=[-0.4,0.4]$. Additionally, we choose masses within $\log(M_{\star}/M_{\odot})=[10.0,11.4]$, and sizes in the range $r_{\mathrm{eff}}\,\mathrm{(arcsec)}=[0.2,1.0]$. The criteria are designed to represent the bulk of typical star-forming galaxies over a mass range large enough to explore trends with stellar mass.
The quantities used in this selection are derived from the 3D-HST catalogue by Momcheva et al. (2016) (see also Brammer et al., 2012; Skelton et al., 2014), which serves as the parent sample for this study and global galaxy properties such as stellar mass and star formation rates are derived following the prescriptions by Wuyts et al. (2011a). Figure 1 provides a visualization of the sample distribution in the SFR vs $M_{\star}$ plane across the two redshift bins.
@@
Figure 2: Number of galaxies observed in each filter. The most frequently observed filter combination—F090W, F115W, F150W, F200W, F277W, F356W, F410M, and F [^4] W—is shared by 739 galaxies (51% of the sample), meaning these galaxies were observed in all these filters. The orange dots indicate the empirical PSF FWHM in each filter. To reduce overcrowding, we omit F250M and F300M, which contain only 10 galaxies each.
For each galaxy, we use a uniform pixel scale of 0.025 arcsec per pixel and retrieve the NIRCam images across all available filters covering the range $0.9\,\mu\mathrm{m}<\lambda<4.8\,\mu\mathrm{m}$. Among the used filters, 7 are broadband (F090W, F115W, F150W, F200W, F277W, F356W, and F [^4] W), and 9 are mediumband (F182M, F210M, F250M, F300M, F335M, F410M, F430M, F460M, and F480M). The most commonly available filter combination in our dataset consists of eight filters—F090W, F115W, F150W, F200W, F277W, F356W, F410M, and F [^4] W— and covers 739 galaxies, making it particularly useful for assessing filter dependence in morphological measurements. Our sample consists of a total of 1,451 galaxies (9877 NIRCam images in total), with 811 and 640 in the low- and high-redshift bins, respectively. Figure 2 presents the number of galaxies observed in each filter along with the corresponding empirical PSF size in arcseconds [^1] Empirical PSF sizes used in this work are taken from https://jwst-docs.stsci.edu/jwst-near-infrared-camera. As shown by Lee et al. (2024), using empirically derived PSFs from field stars in different deep fields has a negligible impact on the inferred morphological fractions across all fields and introduces only minor systematic biases in Sérsic indices.
@@
Figure 3: Histograms showing the distribution of rotational velocity $v_{\mathrm{rot}}(r_{\mathrm{eff}})$ (left) and velocity dispersion $\sigma_{0}$ (right) for different galaxy samples including the RC100 (Nestor Shachar et al., 2023), and those from Übler et al. (2019), Wuyts et al. (2016), Wisnioski et al. (2019), and Price et al. (2016). The legend indicates the datasets included, with the number of galaxies in each sample shown in parentheses. The “All” sample (blue) represents the full dataset, while other colours correspond to specific subsamples from various studies.
@@ 2.2 Reduction of imaging data
We evaluated two distinct data reduction approaches: the public releases from the DAWN JWST Archive [^2] https://dawn-cph.github.io/dja/ reduction (v7.0 + v7.1) and JADES [^3] https://archive.stsci.edu/hlsp/jades (Eisenstein et al., 2023) (v2.0) as well as our custom reduction pipeline. The DAWN Archive provides high-quality reductions of JWST NIRCam data using the grizli reduction framework (Brammer, 2025), which includes baseline calibrations, astrometric alignment using DrizzlePac, and background subtraction optimized for faint structure preservation and noise minimization. The JADES reduction products, while also high-quality, place greater emphasis on precise background modelling and outlier rejection via median mosaics, whereas Grizli (DAWN) applies global sky corrections and sigma-clipping during stacking. For astrometric alignment, JADES employs Montage3, which offers precise registration across multiple exposures, differing from the DrizzlePac-based approach used in DAWN.
@@
In parallel, we developed a custom reduction pipeline tailored to the specific requirements of our analysis. The reduction follows the standard JWST image processing stages [^4] https://jwst-docs.stsci.edu/jwst-science-calibration-pipeline/stages-of-jwst-data-processing: i) converting detector volt signals into rates of units of data numbers per second (DN/s); ii) calibrating rates into flux densities and assigning world coordinates; and iii) creating mosaic images from individual exposures while median-clipping outlier pixels. Building on previous studies (e.g., Bagley et al. 2023; M. Franco et al. in prep.), we also: i) improved removal of NIRCam imaging $1/f$ noise by detecting and subtracting horizontal and vertical stripes; ii) enhanced removal of “snowballs” and “claws” through a combination of manual masking and template fitting, using the NIRCam team’s templates [^5] https://jwst-docs.stsci.edu/depreciated-jdox-articles/nircam-claws-and-wisps. The latest templates were released recently (https://www.stsci.edu/contents/news/jwst/2024/new-nircam-wisp-templates-are-now-available), but as our visually inspected targets are unaffected by claws, we did not update our reduction with the latest templates.; iii) improved world coordinate alignment (astrometry correction) by first detecting bright unsaturated stars/galaxies using sextractor (Bertin & Arnouts, 1996) and then cross-matching them with 3D-HST reference galaxy catalogues, ensuring alignment of JWST images with 3D-HST images; iv) additional flat background subtraction at the end of stage 2 using the Python photutils package before mosaicking to homogenize the mosaic background.
@@
Figure 4: Classification scheme used in the visual classification, with examples from the galaxy sample analyzed in this study. The green, yellow, purple and red coloured boxes indicate the four main categories: Disk, Compact/Spheroidal, Interacting/Merger, and Undefined. The “+ Companion” and “+ Lens” classifications are non-mutually exclusive, meaning a galaxy could belong to any of the main categories while also being identified to have a companion or a lensed system. The blue squares inside the disk category correspond to the five non-exclusive morphological features that a disk galaxy can have.
@@
All scripts and manual mask regions are publicly available on GitHub [^6] https://github.com/1054/Crab.Toolkit.JWST. We processed the imaging data as soon as they became available in the JWST MAST archive from various programs, frequently updating the Calibration Reference Data System (CRDS) and JWST pipeline versions. For this study, the final products used for the morphology analysis were processed using CRDS calibration versions 1023–1130 and the JWST pipeline versions 1.9.0–1.11.4. The final images were produced with a pixel scale of 25 mas, aligned with North up, to facilitate the morphological analysis.
@@
Figure 5: Example colour composite images of galaxies classified as barred (top row) and spiral (bottom row). These categories are not mutually exclusive, as many barred galaxies also exhibit clear spiral arms. The filters used to create each composite image are labelled in the top left corner of each panel. The numbers in the bottom row indicate the number of spiral arms identified in the visual classification. The full mosaics of the barred and spiral galaxies can be found in the following two links: bars and spirals.

### v2

@@
11institutetext: Max-Planck-Institut für extraterrestische Physik (MPE), Giessenbachstraße 1., 85748 Garching, Germany 22institutetext: Max-Planck-Institut für Astrophysik (MPA), Karl-Schwarzschild-Str. 1, 85748, Garching, Germany 33institutetext: Department of Physics, University of Oxford, Clarendon Laboratory, Parks Road, OX1 3PU Oxford, UK 44institutetext: Purple Mountain Observatory, Chinese Academy of Sciences, 10 Yuanhua Road, Nanjing 210023, China 55institutetext: Department of Physics, University of Bath, Claverton Down, Bath BA2 7AY, UK 66institutetext: Research School of Astronomy and Astrophysics, Australian National University, Weston Creek, ACT 2611, Australia 77institutetext: ARC Centre of Excellence for All-Sky Astrophysics in 3 Dimensions (ASTRO 3D), Canberra, ACT 2611, Australia 88institutetext: Kavli Institute for Astronomy and Astrophysics, Peking University, Beijing 100871, People’s Republic of China 99institutetext: Cosmic Dawn Center (DAWN), Denmark 1010institutetext: DTU-Space, Technical University of Denmark, Elektrovej 327, DK2800 Kgs. Lyngby, Denmark $1$institutetext: Department of Physics and Astronomy and PITT PACC, University of Pittsburgh, Pittsburgh, PA 15260, USA 1212institutetext: Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, Padova I-35122, Italy 1313institutetext: School of Physics and Astronomy, Tel Aviv University, Tel Aviv 69978, Israel 1414institutetext: Universidad de Concepción, Víctor Lamas 1290, Barrio Universitario, Concepción, Chile
@@
Juan M. Espejo Salcedo 11 jespejo@mpe.mpg.de    Stavros Pastras 1122    Josef Vácha 33    Claudia Pulsoni 11    Reinhard Genzel 11    N. M. Förster Schreiber 11    Jean-Baptiste Jolly 11    Capucine Barfety 11    Jianhang Chen 11    Giulia Tozzi 11    Daizhong Liu 1144    Lilian L. Lee 11    Stijn Wuyts 55    Linda J. Tacconi 11    Ric Davies 11    Hannah Übler 11    Dieter Lutz 11    Emily Wisnioski 6677    Jinyi Shangguan 1188    Minju Lee 991010    Sedona H. Price $1$    Frank Eisenhauer 11    Alvio Renzini 1212    Amit Nestor Shachar 1313    Rodrigo Herrera-Camus 1414
(Received 24 March 2025 / Accepted 11 June 2025)
The way in which radial flows shape galaxy structure and evolution remains an open question. Internal drivers of such flows, such as bars and spiral arms, known to mediate gas flows in the local Universe, are now observable at high redshift thanks to JWST’s unobscured view. We investigated the morphology of massive star-forming galaxies at $1.30.8<z<$ 1.3 and $2.52.0<z<$ 2.5, epochs marking the peak and decline of cosmic star formation, both well covered by kinematic surveys. Using JWST/NIRCam imaging, we visually classified 1,451 galaxies, identified non-axisymmetric features, counted the number of spiral arms, analyzed nonparametric morphological indicators, and studied the dynamical support of the sample covered by kinematics ($\approx 11\%$ of the sample) as measured via $v/\sigma$. Disk galaxies dominate the sample (fraction $0.82\pm 0.03$); among them, $0.48\pm 0.04$ exhibit spiral structure and $0.11\pm 0.03$ host bars. Both fractions decline with redshift, in agreement with previous studies. The proportion of two- and three-armed spirals remains largely unchanged across our redshift bins: approximately two-thirds show two arms and one-third show three arms in both bins. Notably, we find a higher incidence of three-armed spirals ($\approx 0.30$) than reported in the local Universe ($\approx 0.20$), suggesting a mild evolution in spiral arm multiplicity. Nonparametric morphological metrics strongly correlate with stellar mass but show no significant redshift evolution. Finally, kinematic analysis reveals a correlation between disk morphology and rotational support: most disks exhibit $v/\sigma>3$ and median values of $v/\sigma>7$ for spirals and $v/\sigma>5$ for barred galaxies. This study establishes a population-wide framework for linking galaxy morphology and dynamics at cosmic noon, providing a key reference for future studies on the role of detailed structural features in galaxy evolution.
@@ 1 Introduction
A fundamental question in galaxy evolution is how early star-forming galaxies assembled the well-ordered structures seen in the present-day Universe. While early morphological studies suggested that high-redshift galaxies were highly irregular and dynamically unstable, kinematic surveys have since revealed that disk-like rotation is widespread at cosmic noon ($31<z<$ 3) (e.g., Förster Schreiber et al. 2009; Law et al. 2009; Wisnioski et al. 2015; Stott et al. 2016; Simons et al. 2017; Harrison et al. 2017; Übler et al. 2019). Integral field spectroscopy has consistently shown that a large fraction of star-forming galaxies at these epochs exhibit strong rotational support, often with velocity dispersions lower than expected for turbulent irregular morphologies. These findings challenged the prevailing view based on HST imaging, which primarily identified galaxies with clumpy chaotic structures (e.g., Abraham et al., 1996; Conselice et al., 2000; Jogee et al., 2004; Lotz et al., 2006; Elmegreen et al., 2007; Sheth et al., 2008).
@@
The arrival of JWST has dramatically transformed our ability to study galaxy structure at high redshift, offering a combination of longer wavelengths, increased depth, and unprecedented resolution. Recent studies leveraging JWST NIRCam imaging consistently reveal a high fraction of disk galaxies at and beyond cosmic noon (e.g., Jacobs et al. 2023; Ferreira et al. 2023; Kartaltepe et al. 2023; Liu et al. 2023; Nelson et al. 2023; Huertas-Company et al. 2024; Lee et al. 2024; Tohill et al. 2024; Pandya et al. 2024; Smethurst et al. 2025). Additionally, some studies provide clear evidence of spiral arms and bars (e.g., Costantin et al. 2023; Ferreira et al. 2023; Guo et al. 2023; Kuhn et al. 2024; Le Conte et al. 2024; Xiao et al. 2025; Chugunov et al. 2025; Géron et al. 2025). Thus, many galaxies that once appeared irregular or clumpy in HST imaging are now recognized as well-ordered disks, some of which exhibit prominent non-axisymmetric features.
Beyond their prevalence, these nonsymmetric morphological features are key agents of secular evolution, regulating gas flows and star formation within galaxies. In the local Universe, bars and spiral arms are known to redistribute angular momentum and funnel gas toward galaxy centers, often fueling central star formation and bulge growth. As demonstrated by Yu et al. (2022) with the EDGE-CALIFA survey of local galaxies, spiral arms and bars both efficiently channel gas inward, leading to higher central gas concentrations and elevated star formation rates.
If these mechanisms were already at play at cosmic noon, they could have profound implications for the early buildup of galaxy structure. Recent kinematic studies have uncovered evidence for rapid radial inflows in high-redshift main-sequence disks (e.g., Genzel et al., 2023, Pastras et al., 2025), which suggests that these processes may have been even more pronounced in the past. These findings contradict earlier expectations that non-axisymmetric structures such as spiral arms and bars would be short-lived in gas-rich turbulent environments (e.g., Genzel et al. 2008; Ceverino et al. 2010; Bournaud et al. 2011). Instead, they suggest that such features were not only prevalent, but they also played a fundamental role in shaping galaxy evolution throughout cosmic time. Given these advancements, establishing a direct connection between morphology and kinematics is crucial for understanding how high-redshift disks evolve and whether non-axisymmetric features are the primary drivers of internal gas flows and structural transformation.
@@
Our goal is to establish a robust morphological foundation for interpreting the growing body of kinematic measurements from near-infrared integral field spectroscopy and millimeter and/or micromillimeter observations. Specifically, our sample selection aligns with that of major spectroscopic surveys such as KMOS3D, SINS, KROSS, NOEMA-3D, and PHIBSS, thus enabling the connections between the morphological and kinematic analyses. By providing a comprehensive visual classification of galaxy structure (including disks, bars, and spiral arms), this work serves as a key reference for future studies whose aim is to investigate the connections between morphology, gas transport, and secular evolution. Additionally, we complement our visual classifications with quantitative (nonparametric) morphological metrics, leveraging well-established statistical methods to characterize galaxy structure and to systematically link morphology and dynamics at cosmic noon.
In Sect. 2, we describe the sample selection, followed by a discussion of the imaging data, its corresponding data reduction methods, and the assembly of the kinematic sample used in this study. Section 3 presents the overall morphological properties of the sample, detailing the visual classification scheme, general classification results, and the identification of bars and spiral structures. Section 4 examines the evolution of these morphological features as a function of redshift and compares our findings with previous studies. In Sect. 5, we perform a quantitative analysis of galaxy morphology, exploring nonparametric structural indicators and their relationship to the visual classifications. In Sect. 6 we investigate the connection between morphology and kinematics, and discuss the implications for galaxy evolution. Finally, Sect. 7 summarizes the key findings of this work. Additional details, including discussions of observational limitations, preparations for the morphological analysis, and a description of the methodology used to separate galaxy morphologies based on structural metrics are provided in Appendices A-C.
@@
Figure 1: Star formation rate (SFR) vs stellar mass $M_{\star}$ for the selected sample. The lines represent the Speagle et al. (2014) relations for the mid-points of the redshift bins $z=1.05$ and $z=2.25$.
@@ 2 Sample
Figure 2: Number of galaxies observed in each filter. The most frequently observed filter combination—F090W, F115W, F150W, F200W, F277W, F356W, F410M, and F [^4] W—is shared by 739 galaxies (51% of the sample), meaning these galaxies were observed in all these filters. The orange dots indicate the empirical PSF FWHM in each filter. To reduce overcrowding, we omitted F250M and F300M, which contain only ten galaxies each.
@@ 2.1 NIRCam imaging
To investigate morphological evolution at cosmic noon, we focus on two epochs corresponding to $1.30.8<z<$ 1.3 and $2.52.0<z<$ 2.5. Since star-forming galaxies dominate the population at these epochs, we apply selection criteria targeting massive star-forming galaxies with star formation rates (SFR) within 0.4 dex of the star formation main sequence (MS), as parameterized by Whitaker et al. (2014), i.e., $\log(\mathrm{SFR}/\mathrm{SFR_{MS}})=[-0.4,0.4]$. Additionally, we choose masses within $\log(M_{\star}/M_{\odot})=[10.0,11.4]$, and sizes in the range $r_{\mathrm{eff}}\,\mathrm{(arcsec)}=[0.2,1.0]$. The criteria are designed to represent the bulk of typical star-forming galaxies at high masses.
[…]
@@
For each galaxy, we use a uniform pixel scale of 0.025 arcsec per pixel and retrieve the NIRCam images across all available filters covering the range $0.9\,\mu\mathrm{m}<\lambda<4.8\,\mu\mathrm{m}$. Among the used filters, seven are broadband (F090W, F115W, F150W, F200W, F277W, F356W, and F [^4] W), and nine are mediumband (F182M, F210M, F250M, F300M, F335M, F410M, F430M, F460M, and F480M). The most commonly available filter combination in our dataset consists of eight filters—F090W, F115W, F150W, F200W, F277W, F356W, F410M, and F [^4] W— and covers 739 galaxies, making it particularly useful for assessing filter dependence in morphological measurements. Our sample consists of a total of 1,451 galaxies (9,877 NIRCam images in total), where 811 and 640 are in the low- and high-redshift bins, respectively. Figure 2 presents the number of galaxies observed in each filter along with the corresponding empirical PSF size in arcseconds [^1] Empirical PSF sizes used in this work are taken from https://jwst-docs.stsci.edu/jwst-near-infrared-camera. As shown by Lee et al. (2024), using empirically derived PSFs from field stars in different deep fields has a negligible impact on the inferred morphological fractions across all fields and introduces only minor systematic biases in Sérsic indices.
@@
Figure 3: Histograms showing the distribution of rotational velocity $v_{\mathrm{rot}}(r_{\mathrm{eff}})$ (left) and velocity dispersion $\sigma_{0}$ (right) for different galaxy samples including the RC100 (Nestor Shachar et al., 2023), and those from Übler et al. (2019), Wuyts et al. (2016), Wisnioski et al. (2019), and Price et al. (2016). The legend indicates the corresponding dataset and includes the number of galaxies in each sample (in parentheses). The “All” sample (blue) represents the full dataset; the other colors correspond to specific subsamples from various studies.
@@ 2.2 Reduction of imaging data
We evaluated two distinct data reduction approaches: the public releases from the DAWN JWST Archive [^2] https://dawn-cph.github.io/dja/ reduction (v7.0 + v7.1) and JADES [^3] https://archive.stsci.edu/hlsp/jades (Eisenstein et al., 2023) (v2.0) as well as our custom reduction pipeline. The DAWN Archive provides high-quality reductions of JWST NIRCam data using the grizli reduction framework (Brammer, 2025), which includes baseline calibrations, astrometric alignment using DrizzlePac, and background subtraction optimized for faint structure preservation and noise minimization. The JADES reduction products, while also high-quality, place greater emphasis on precise background modeling and outlier rejection via median mosaics, whereas grizli (DAWN) applies global sky corrections and sigma-clipping during stacking. For astrometric alignment, JADES employs Montage3, which offers precise registration across multiple exposures, differing from the DrizzlePac-based approach used in DAWN.
@@
In parallel, we developed a custom reduction pipeline tailored to the specific requirements of our analysis. The reduction follows the standard JWST image processing stages [^4] https://jwst-docs.stsci.edu/jwst-science-calibration-pipeline/stages-of-jwst-data-processing: i) converting detector volt signals into rates of units of data numbers per second (DN/s); ii) calibrating rates into flux densities and assigning world coordinates; and iii) creating mosaic images from individual exposures while median-clipping outlier pixels. Building on previous studies (e.g., Bagley et al. 2023; M. Franco et al. in prep.), we also: i) improved removal of NIRCam imaging $1/f$ noise by detecting and subtracting horizontal and vertical stripes; ii) enhanced removal of snowballs and claws through a combination of manual masking and template fitting, using the NIRCam team’s templates [^5] https://jwst-docs.stsci.edu/depreciated-jdox-articles/nircam-claws-and-wisps. The latest templates were released recently (https://www.stsci.edu/contents/news/jwst/2024/new-nircam-wisp-templates-are-now-available), but as our visually inspected targets are unaffected by claws, we did not update our reduction with the latest templates.; iii) improved world coordinate alignment (astrometry correction) by first detecting bright unsaturated stars and/or galaxies using sextractor (Bertin & Arnouts, 1996) and then cross-matching them with 3D-HST reference galaxy catalogs, ensuring alignment of JWST images with 3D-HST images; iv) additional flat background subtraction at the end of stage two using the Python photutils package before mosaicking to homogenize the mosaic background.
Figure 4: Example of the interactive tool GalClass (Pastras, 2025) used for our morphological classification. The displayed galaxy, GS4_07410, is shown in the F090W filter. The top row displays the original image, galfit model, and residuals; the bottom row shows the full-size image, the smoothed image, and corresponding residuals. The left sidebar provides metadata and galfit parameters, while the right panel contains the classification interface.
@@
Figure 5: Classification scheme used in the visual classification, including examples from the galaxy sample analyzed in this study. The colored boxes, green, yellow, purple, and red, indicate the four main categories of disk, compact/spheroidal, interacting/merger, and undefined. The + Companion and + Lens classifications are nonmutually exclusive, meaning a galaxy could belong to any of the main categories, while also being identified as having a companion or a lensed system. The blue squares inside the disk category correspond to the five non-exclusive morphological features that a disk galaxy can have.
@@
All scripts and manual mask regions are publicly available on GitHub [^6] https://github.com/1054/Crab.Toolkit.JWST. We processed the imaging data as soon as they became available in the JWST MAST archive from various programs, frequently updating the Calibration Reference Data System (CRDS) and JWST pipeline versions. For this study, the final products used for the morphology analysis were processed using CRDS calibration versions 1023–1130 and the JWST pipeline versions 1.9.0–1.11.4. The final images were produced using a pixel scale of 25 mas, aligned with North up, to facilitate the morphological analysis.
@@
Figure 6: Example color composite images of galaxies classified as barred (top row) and spiral (bottom row). These categories are nonmutually exclusive, as many barred galaxies also exhibit clear spiral arms. The filters used to create each composite image are labeled in the top left corner of each panel. The numbers in the bottom row indicate the number of spiral arms identified in the visual classification. The full mosaics of the barred and spiral galaxies can be found at the following two links: bars and spirals.